# Computing Chebyshev polynomials using the complex Remez algorithm

### v2

@@ 4.3.2 Circular lune
Based on the examples of the regular polygons and circular lunes together with our belief that Conjecture 2 is valid we conjecture the following result which is a partial reformulation of [29, Theorem 1.5] to the setting of Chebyshev polynomials. We define a singularity point of a piecewise analytic curve as a point where the derivative of the arc-length parametrization of the curve has limits from either sides but form an angle $\lambda\neq\pi$ to each other.
Conjecture 3.
Let $\mathsf{E}\subset\mathbb{C}$ denote the closure of a Jordan domain with piecewise analytic boundary such that $\partial\mathsf{E}$ has a singularity other than an outward cusp. Then there is a subsequence $\{T_{n_{k}}^{\mathsf{E}}\}$ such that
$\nu(T_{n_{k}}^{\mathsf{E}})\xrightarrow{\ast}\mu_{\mathsf{E}},\quad n_{k}\rightarrow\infty.$
4.3.3 Hypocycloid
The reason that an outward cusp is excluded in Conjecture 3 is that the result does not hold in the Faber setting if the bounding curve has an outward cusp as is shown in [24]. Indeed, exactly as is the case for Faber polynomials, we believe that the example of an hypocycloid provides an example where the zeros of $T_{n}^{\mathsf{E}}$ do not approach all of the boundary. It is clearly suggested by Figures 5(g)-5(j) that the support of $\nu(T_{n}^{\mathsf{H}_{m}})$ is confined to the diagonals between the cusps and the origin for all values of $n$ computed. This is in accordance with the behavior exhibited by $\nu(F_{n}^{\mathsf{H}_{m}})$ and we believe that an analogous result as [24, Theorem 3.1] is true in this case.
Conjecture 4.
The zeros of $T_{n}^{\mathsf{H}_{m}^{r}}$ for $r\geq 1$ are confined to the set
$\left\{te^{2\pi ik/m}:0\leq t\leq\frac{m}{m-1},\,k=0,1,\dotsc,m-1\right\}.$ (25)
Again, if we choose to believe Conjecture 2 then this together with Conjecture 25 would imply that the zeros of $T_{n}^{\mathsf{H}_{m}^{r}}$ would move along the straight diagonals as $r$ increases and approach the corresponding zeros of the Faber polynomials, this is something we find reasonable to believe.
On the other hand, we note that numerical simulations indicate that the zeros of the corresponding Bergman polynomials corresponding to $\mathsf{H}_{m}$ and its interior, all lie on the straight lines in (25) for small degrees. However it follows from [38, Theorem 2.1] that at least a subsequence of the Bergman polynomials have zero counting measures converging weak-star to $\mu_{\mathsf{H}_{m}}$.
4.3.4 Lemniscate
Recall that $\mathsf{L}^{r}_{m}=\{z:|z^{m}-1|=r^{m}\}$, $\mathsf{L}^{r}_{2}=\mathsf{L}^{r}$ and that $\mathsf{L}=\mathsf{L}^{1}$. Based on Figure 5(k) it seems reasonable to assume that
$\lim_{n\rightarrow\infty}\nu(T_{2n+1}^{\mathsf{L}})(M)=0$ (26)
@@ 4.3.4 Lemniscate
Corresponding to each weight of the form $|\zeta+1|^{s}$ for $s\geq 0$ there is a minimizing weighted Chebyshev polynomial which we denote with $T_{n}^{s}$, see [5]. In the particular case where $s=1$ it is shown in [5, Theorem 3] that $\nu(T_{n}^{1})$ converges weak-star to equilibrium measure on $\mathbb{T}$. This implies that an analogous result as (26) is valid for compact subsets of $\mathbb{D}$. There is no reason to believe that such a result should exclusively hold for the parameter value of $s=1$ and we therefore suspect that
@@
We turn our attention to the outer lemniscates $\mathsf{L}^{r}$ with $r>1$. Surprisingly, based on Figure 5(l) it seems like the zeros of $T_{2n+1}^{\mathsf{L}^{r}}$ all lie strictly inside $\mathsf{L}$ except for the single zero at 00. Although the main Theorem in [40] implies that the zeros asymptotically stay away from $\mathsf{L}^{r}$ there is no results hinting toward the fact the zeros seem to cluster on $\mathsf{L}$. If one believes Conjecture 2 so that $T_{n}^{\mathsf{L}^{r}}\rightarrow F_{n}^{\mathsf{L}}$ as $r\rightarrow\infty$ then it is reasonable to assume that the zeros of $T_{2n+1}^{\mathsf{L}^{r}}$ lie on or inside $\mathsf{L}$ for all values of $n$ and $r$ since the zeros of $F_{2n+1}^{\mathsf{L}}$ have this very behavior.
Analogous results seem to hold true with $\mathsf{L}^{r}$ replaced by $\mathsf{L}_{m}^{r}$ for any value of $m$ as the corresponding numerical simulations indicate the same pattern. Generalizations of Ullmans result concerning the asymptotic zero distribution of the Faber polynomials on $\mathsf{L}_{m}$ can be found in [21]. We further believe that a general version of the above result can be formulated for any connected lemniscate. To understand this perspective we introduce the notion of a critical value of a polynomial. This is a number $P(z)$ where $z$ is such that $P^{\prime}(z)=0$. The polynomial $z^{2}-1$ has one critical value, namely $-1$ which is attained at the origin. This implies that the curve $\mathsf{L}=\{z:|z^{2}-1|=1\}$ will contain a critical point of $z^{2}-1$ resulting in the fact that the curve forms a crossing with itself at the origin. In general, if $c$ is a critical value of a polynomial $P$ then $\{z:|P(z)|=|c|\}$ will contain a crossing point.
If we consider the polynomial $Q(z)=z^{4}-z^{2}$ then $Q$ has two critical values, namely $1/4$ and 00. Upon inspection of Figure 5(m) it becomes apparent that the zeros of the Chebyshev polynomials on the curve $\{z:|z^{4}-z^{2}|=5/4\}$ seem to approach the critical curve $\{z:|z^{4}-z^{2}|=1/4\}$ which correspond to the lemniscate where the largest critical value is attained (in modulus). Equivalently, this curve is characterized by being the curve $\{z:|z^{4}-z^{2}|=r\}$ with smallest value of $r>0$ which is connected.
A similar pattern emerges for the lemniscates of the form $\{z:|z^{3}+z+1|=r\}$ with $r\geq\sqrt{31/27}$, see Figure 5(n). For the polynomial $P(z)=z^{3}+z+1$ the critical point is ±$\pm i/\sqrt{3}$ and the corresponding critical value is $1\pm i2/3\sqrt{3}$. Since $|1+i2/3\sqrt{3}|=\sqrt{31/27}$ we see that the critical lemniscate corresponds to $r=\sqrt{31/27}$. Again, this critical lemniscate seems to attract the zeros of the Chebyshev polynomials corresponding to larger values of $r$. We believe that this can be formulated as a general result as we have observed this very behavior for all lemniscates that we have considered.
Conjecture 5.
Let $P$ be a polynomial of degree $m$ with largest critical value in terms of absolute value given by $c$. For any $r\geq|c|$ let
@@
then for a fixed $l\in\{1,\dotsc,m-1\}$
@@
as $n\rightarrow\infty$.
Based on Figures 5(m) and 5(n) this seems to be the case. Observe that (5) implies that $T_{nm}^{\mathsf{E}^{r}}=a^{-n}P(z)^{n}$ where $a$ is the leading coefficient of $P$ in which case the zero counting measure is constant.
It could be further speculated what happens in the general case for level curves of conformal maps. Assume that $\mathsf{E}$ is a connected compact set with simply connected complement and $\Phi:\mathbb{C}\setminus\mathsf{E}\rightarrow\{z:|z|>1\}$ is the conformal map of the form $\Phi(z)=\operatorname{Cap}(\mathsf{E})^{-1}z+O(1)$ as $z\rightarrow\infty$. Again, introducing the set $\mathsf{E}^{r}=\{z:|\Phi(z)|=r\}$ then the bounding curve of $\mathsf{E}^{r}$ is analytic for $r>1$. From [40] we know that the zeros of $T_{n}^{\mathsf{E}^{r}}$ asymptotically stay away from the boundary, in the sense that there exists a neighborhood of the boundary where $T_{n}^{\mathsf{E}^{r}}$ is zero free for large $n$. The question is if something similar as in the case of lemniscates happens in this situation. Do the zeros asymptotically approach $\mathsf{E}$? This is true for the corresponding Faber polynomials and therefore the validity of Conjecture 2 could hint at this being true for the corresponding Chebyshev polynomials.
4.4 Concluding remarks
@@
Based on the fact that $T_{2n}^{\mathsf{L}}(z)=(z^{2}-1)^{n}$ it was suggested in [10] that the odd Chebyshev polynomials $T_{2n+1}^{\mathsf{L}}$ which apart from having a zero at the origin should behave similarly. Explicitly it is written on [10, p. 215] that “…we suspect (but cannot prove) that for $j$ large all the other zeros of $T_{2j+1}$ lie in small neighborhoods of ±$\pm 1$ and that the above $d\mu_{\infty}$ is also the limit through odd $n$’s.” Here $d\mu_{\infty}=\frac{1}{2}(\delta_{-1}+\delta_{1})$. We initially set out to show this. Since we did not progress in this regard we started considering numerical methods to compute the Chebyshev polynomials with the intent of better understanding how the zeros approached ±$\pm 1$. Using Tang’s algorithm we could compute the Chebyshev polynomials corresponding to $\mathsf{L}$ and the result surprised us. The zeros seemed to behave opposite to our conjecture and approached the bounding curve rather than the two points ±$\pm 1$. The use of the algorithm therefore showed us that the hypothesis we initially had believed was probably incorrect and that our conjecture should be modified. We made partial progress in proving Conjecture 3 in [5] by showing that a related problem satisfied the conjectured behavior. However, we are still lacking a complete proof of this.
With the algorithm at hand we set out to study Chebyshev polynomials corresponding to a wide variety of sets whose asymptotic behavior remain unknown. We believe that making use of the algorithm is a good way of getting predictions on the behavior of Chebyshev polynomials. The results in [5] and [8] are based on conjectures formulated using initial numerical experiments. Some rather surprising results have also been suggested to us by numerical experiments along the way. In particular, the relation between Faber polynomials and Chebyshev polynomials specified in Conjecture 2 does not seem to have been given any attention in the literature in the past although the fact that they coincide for certain sets is known.
In short, we believe that use of Tang’s algorithm in the study of Chebyshev polynomials may prove useful in the future when formulating conjectures on their asymptotic behavior.
@@ Appendix A Tang’s algorithm
$L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(g)=\sum_{j=1}^{n+1}r_{j}\mathrm{Re}(e^{-i\alpha_{j}}g(z_{j}))$ (27)
@@
$A(\boldsymbol{z},\boldsymbol{\alpha})=\begin{pmatrix}1&1&\cdots&1\\ \operatorname{Re}(e^{-i\alpha_{1}}\varphi_{1}(z_{1}))&\operatorname{Re}(e^{-i\alpha_{2}}\varphi_{1}(z_{2}))&\cdots&\operatorname{Re}(e^{-i\alpha_{n+1}}\varphi_{1}(z_{n+1}))\\ \vdots&\vdots&\ddots&\vdots\\ \operatorname{Re}(e^{-i\alpha_{1}}\varphi_{n}(z_{1}))&\operatorname{Re}(e^{-i\alpha_{2}}\varphi_{n}(z_{2}))&\cdots&\operatorname{Re}(e^{-i\alpha_{n+1}}\varphi_{n}(z_{n+1}))\end{pmatrix}$ (28)
@@
$c_{f}(\boldsymbol{z},\boldsymbol{\alpha})=\begin{pmatrix}\operatorname{Re}(e^{-i\alpha_{1}}f(z_{1}))\\ \vdots\\ \operatorname{Re}(e^{-i\alpha_{n+1}}f(z_{n+1}))\end{pmatrix}.$ (29)
@@
$L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(f)=c_{f}(\boldsymbol{z},\boldsymbol{\alpha})^{T}\boldsymbol{r}$ (30)
@@
$A(\boldsymbol{z},\boldsymbol{\alpha})\boldsymbol{r}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}.$ (31)
Parameters $\boldsymbol{r},\,\boldsymbol{\alpha},\,\boldsymbol{z}$ satisfying (31) are called admisible if additionally $A(\boldsymbol{z},\boldsymbol{\alpha})$ is invertible. If $\varphi^{\ast}=\sum_{k=1}^{n}\lambda_{k}^{\ast}\varphi_{k}$, $\lambda_{k}\in\mathbb{R}$, is a best approximation and $\boldsymbol{r}^{\ast},$ $\boldsymbol{\alpha}^{\ast}$, $\boldsymbol{z}^{\ast}$ are corresponding admissible parameters such that $L_{\boldsymbol{r}^{\ast},\boldsymbol{\alpha}^{\ast},\boldsymbol{z}^{\ast}}(f)=\|f-\varphi^{\ast}\|_{\mathsf{E}}$ then
$A(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})^{T}\begin{pmatrix}\|f-\varphi^{\ast}\|_{\mathsf{E}}\\ \lambda_{1}^{\ast}\\ \vdots\\ \lambda_{n}^{\ast}\end{pmatrix}=c_{f}(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$ (32)
and therefore if $A(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$ is invertible we can recover the extremal coefficients $\lambda_{1}^{\ast},\dotsc,\lambda_{n}^{\ast}$ from $A(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$ and $c_{f}(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$. We assume, as in [45, 46, 16] that $\mathsf{E}=I=:[0,1]$ (which is no restriction) since we can always parametrize $\mathsf{E}$ using $[0,1]$. To emphasize that we are working on $[0,1]$ we let $\boldsymbol{z}=\boldsymbol{t}=\{t_{j}\}$.
For an implementation of the algorithm in Python, see [37].
1
Input: admissible parameters $\{t_{j}\}_{j=1}^{n+1}$, $\{\alpha_{j}\}_{j=1}^{n+1}$, $\{r_{j}\}_{j=1}^{n+1}$
compute: $\begin{bmatrix}\ast&\boldsymbol{\lambda}\end{bmatrix}^{T}=A(\boldsymbol{t},\boldsymbol{\alpha})^{-1}c_{f}(\boldsymbol{t},\boldsymbol{\alpha})$
@@
let: $\varphi=\sum_{k=1}^{n}\lambda_{k}\varphi_{k}$
2
3if $\|f-\varphi\|_{I}=c_{f}(\boldsymbol{t},\boldsymbol{\alpha})^{T}\boldsymbol{r}$ then
@@
       let: $\boldsymbol{d}$ be a solution to $\boldsymbol{d}\cdot A(\boldsymbol{t},\boldsymbol{\alpha})^{T}=\begin{bmatrix}1&\operatorname{Re}(e^{-i\vartheta}\varphi_{1}(x))&\cdots&\operatorname{Re}(e^{-i\vartheta}\varphi_{n}(x))\end{bmatrix}$
@@
       let: $t_{\rho}:=x$
       let: $\alpha_{\rho}:=\vartheta$
7       for $j\in\{1\cdots n+1\}\setminus\{\rho\}$ do
8             $r_{j}:=r_{j}-\delta\cdot d_{j}$
9      $r_{\rho}=\delta$
10
Algorithm 1 Remez Exchange
1
Input: Basis functions $\{\varphi_{k}\}_{k=1}^{n}$, $f\in\mathcal{C}_{\mathbb{R}}(I)$, $\{t_{j}^{(1)}\}_{j=1}^{n+1}$, $\{\alpha_{j}^{(1)}\}_{j=1}^{n+1}$
find: $\boldsymbol{r}^{(1)}$ satisfying $A(\boldsymbol{t}^{(1)},\boldsymbol{\alpha}^{(1)})\boldsymbol{r}^{(1)}=\begin{bmatrix}1\\ \boldsymbol{0}\end{bmatrix}$.
2 for $j\in\{1,\dotsc,n+1\}$ do
3       if $r_{j}^{(1)}<0$ then
             let: $\alpha_{j}^{(1)}:=\alpha_{j}^{(1)}\pm\pi\in[0,2\pi)$
4
5
  // This will produce admissible parameters
6 for $\nu=1,2,\cdots$ do
       let: $\boldsymbol{t}^{(\nu+1)},\boldsymbol{\alpha}^{(\nu+1)},\boldsymbol{r}^{(\nu+1)},\boldsymbol{\lambda}^{(\nu+1)}$ := RemezExchange($\boldsymbol{t}^{(\nu)},\boldsymbol{\alpha}^{(\nu)},\boldsymbol{r}^{(\nu)}$)
7
      let: $\varphi^{(\nu+1)}:=\sum_{k=1}^{n}\lambda_{k}^{(\nu+1)}\varphi_{k}$.
8       if $\|f-\varphi^{(\nu+1)}\|_{I}-L_{\boldsymbol{t}^{(\nu+1)},\boldsymbol{\alpha}^{(\nu+1)},\boldsymbol{r}^{(\nu+1)}}(f)<\mathrm{threshold}$ then
             return: $\boldsymbol{\lambda}^{\nu+1}$
              // Sufficient approximation $\varphi^{(\nu+1)}$ found.
9
10
11
@@ 9      $r_{\rho}=\delta$
Algorithms 1 and 2 enable the computation of best approximations to complex valued functions.
Acknowledgement
The author is grateful to Frank Wikström and Jacob Stordal Christiansen, both at Lund University. The former for providing excellent advice on the presentation of the numerical experiments in this manuscript and the latter for helpful discussions on the material presented herein.

### v3

@@ 4.3.2. Circular lunes
4.3.3. Hypocycloids
If the bounding curve has an outward cusp, then the example of Hypocycloids illustrate that the zeros of Faber polynomials need not approach all of the boundary, as shown in [26]. We believe that the zeros of the corresponding Chebyshev polynomials behave analogously. It is clearly suggested by Figures 7(a)-7(d) that the support of $\nu(T_{n}^{\mathsf{H}_{m}})$ is confined to the diagonals between the cusps and the origin. This is in accordance with the behavior exhibited by $\nu(F_{n}^{\mathsf{H}_{m}})$ and we believe that an analogous result as [26, Theorem 3.1] is true in this case.
That is to say, we believe that the zeros of $T_{n}^{\mathsf{H}_{m}^{r}}$ for $r\geq 1$ are confined to the set
(31) $\left\{te^{2\pi ik/m}:0\leq t\leq\frac{m}{m-1},\,k=0,1,\dotsc,m-1\right\}.$
If we choose to believe the connection between Chebyshev polynomials and Faber polynomials on equipotential lines from (30) then this would imply that the zeros of $T_{n}^{\mathsf{H}_{m}^{r}}$ would move along the diagonals (31) as $r$ increases and approach the corresponding zeros of the Faber polynomials, this is something we find reasonable to believe.
On the other hand, we note that numerical simulations indicate that the zeros of the corresponding Bergman polynomials corresponding to $\mathsf{H}_{m}$ and its interior, all lie on the straight lines in (31) for small degrees. However, it follows from [43, Theorem 2.1] that at least a subsequence of the Bergman polynomials have zero counting measures converging weak-star to $\mu_{\mathsf{H}_{m}}$.
4.3.4. Lemniscates
Recall that $\mathsf{L}^{r}_{m}=\{z:|z^{m}-1|=r^{m}\}$, $\mathsf{L}^{r}_{2}=\mathsf{L}^{r}$ and that $\mathsf{L}=\mathsf{L}^{1}$. Based on Figure 8(a) it seems reasonable to assume that
(32) $\lim_{n\rightarrow\infty}\nu(T_{2n+1}^{\mathsf{L}})(M)=0$
@@ 4.3.4. Lemniscates
Corresponding to each weight of the form $|\zeta+1|^{s}$ for $s\geq 0$ there is a minimizing weighted Chebyshev polynomial which we denote with $T_{n}^{s}$, see [6]. In the particular case where $s=1$ it is shown in [6, Theorem 3] that $\nu(T_{n}^{1})$ converges in the weak-star sense to the equilibrium measure on $\mathbb{T}$. This implies that an analogous result as (32) is valid for compact subsets of $\mathbb{D}$. There is no reason to believe that such a result should exclusively hold for the parameter value of $s=1$ and we therefore suspect that
@@
We turn our attention to the outer lemniscates $\mathsf{L}^{r}$ with $r>1$. Surprisingly, based on Figure 8(b) it seems like the zeros of $T_{2n+1}^{\mathsf{L}^{r}}$ all lie strictly inside $\mathsf{L}$ except for the single zero at 00. Although the main Theorem in [45] implies that the zeros asymptotically stay away from $\mathsf{L}^{r}$, there is no results hinting toward the fact the zeros seem to cluster on $\mathsf{L}$. If one believes in (30), as argued for in Section 4.2, we find it reasonable to believe that the zeros of $T_{2n+1}^{\mathsf{L}^{r}}$ lie on or inside $\mathsf{L}$ for all values of $n$ and $r$ since the zeros of $F_{2n+1}^{\mathsf{L}}$ have this very behavior.
Analogous results seem to hold true with $\mathsf{L}^{r}$ replaced by $\mathsf{L}_{m}^{r}$ for any value of $m$ as the corresponding numerical simulations indicate the same pattern. Generalizations of Ullmans result concerning the asymptotic zero distribution of the Faber polynomials on $\mathsf{L}_{m}$ can be found in [23].
We further believe that a similar picture holds for any connected lemniscate. To make precise our belief, we introduce the notion of a critical value of a polynomial. This is a number $P(z)$ where $z$ is such that $P^{\prime}(z)=0$. The polynomial $z^{2}-1$ has one critical value, namely $-1$ which is attained at the origin. This implies that the curve $\mathsf{L}=\{z:|z^{2}-1|=1\}$ will contain a critical point of $z^{2}-1$ resulting in the fact that the curve forms a crossing with itself at the origin. In general, if $c$ is a critical value of a polynomial $P$ then $\{z:|P(z)|=|c|\}$ will contain a crossing point.
If we consider the polynomial $Q(z)=z^{4}-z^{2}$ then $Q$ has two critical values, namely $1/4$ and 00. Upon inspection of Figure 8(c) it becomes apparent that the zeros of the Chebyshev polynomials on the curve $\{z:|z^{4}-z^{2}|=5/4\}$ seem to approach the critical curve $\{z:|z^{4}-z^{2}|=1/4\}$ which correspond to the lemniscate where the largest critical value is attained (in modulus). Equivalently, this curve is characterized by being the curve $\{z:|z^{4}-z^{2}|=r\}$ with smallest value of $r>0$ which is connected.
A similar pattern emerges for the lemniscates of the form $\{z:|z^{3}+z+1|=r\}$ with $r\geq\sqrt{31/27}$, see Figure 8(d). For the polynomial $P(z)=z^{3}+z+1$ the critical point is ±$\pm i/\sqrt{3}$ and the corresponding critical value is $1\pm i2/3\sqrt{3}$. Since $|1+i2/3\sqrt{3}|=\sqrt{31/27}$ we see that the critical lemniscate corresponds to $r=\sqrt{31/27}$. Again, this critical lemniscate seems to attract the zeros of the Chebyshev polynomials corresponding to larger values of $r$.
We question if this holds in greater generality. Let $P$ be a polynomial of degree $m$ with largest critical value in terms of absolute value given by $c$. For any $r\geq|c|$ let
@@
then for a fixed $l\in\{1,\dotsc,m-1\}$ is it true that
@@
as $n\rightarrow\infty$?
Based on Figures 8(c) and 8(d) this seems to be the case. Observe that (5) implies that $T_{nm}^{\mathsf{E}^{r}}=a^{-n}P(z)^{n}$ where $a$ is the leading coefficient of $P$ in which case the zero counting measure is constant.
It could be further speculated what happens in the general case for level curves of conformal maps. Assume that $\mathsf{E}$ is a connected compact set with simply connected complement and $\Phi:\mathbb{C}\setminus\mathsf{E}\rightarrow\{z:|z|>1\}$ is the conformal map of the form $\Phi(z)=\operatorname{Cap}(\mathsf{E})^{-1}z+O(1)$ as $z\rightarrow\infty$. Again, introducing the set $\mathsf{E}^{r}=\{z:|\Phi(z)|=r\}$ then the bounding curve of $\mathsf{E}^{r}$ is analytic for $r>1$. From [45] we know that the zeros of $T_{n}^{\mathsf{E}^{r}}$ asymptotically stay away from the boundary, in the sense that there exists a neighborhood of the boundary where $T_{n}^{\mathsf{E}^{r}}$ is zero free for large $n$. The question is if something similar as in the case of lemniscates happens in this situation. Do the zeros asymptotically approach $\mathsf{E}$? This is true for the corresponding Faber polynomials which could hint at this being true for the corresponding Chebyshev polynomials.
4.4. Concluding remarks
@@
Based on the fact that $T_{2n}^{\mathsf{L}}(z)=(z^{2}-1)^{n}$ it was suggested in [11] that the odd Chebyshev polynomials $T_{2n+1}^{\mathsf{L}}$, which apart from having a zero at the origin, should behave similarly. Explicitly it is written in [11, p. 215] that
“…we suspect (but cannot prove) that for $j$ large all the other zeros of $T_{2j+1}$ lie in small neighborhoods of ±$\pm 1$ and that the above $d\mu_{\infty}$ is also the limit through odd $n$’s.”
Here $d\mu_{\infty}=\frac{1}{2}(\delta_{-1}+\delta_{1})$. We initially set out to show this. Since we did not progress in this regard we started considering numerical methods to compute the Chebyshev polynomials with the intent to better understand how the zeros approached ±$\pm 1$. Using Tang’s algorithm we could compute the Chebyshev polynomials corresponding to $\mathsf{L}$ and the result surprised us. The zeros seemed to behave opposite to what was conjectured. They approached the bounding curve rather than the two points ±$\pm 1$. The use of Tang’s algorithm therefore showed us that the hypothesis we initially had believed was probably incorrect and that the conjecture should be modified. Partial progress to understanding the zeros of Chebyshev polynomials relative to lemniscates was done in [6] by studying a related problem. However, we are still lacking a complete proof of this.
With Tang’s algorithm at hand, we set out to study Chebyshev polynomials corresponding to a wide variety of sets whose asymptotic behavior remain unknown. We believe that making use of Tang’s algorithm (or other related algorithms such as the Lawson algorithm) is a good way of getting predictions on the behavior of Chebyshev polynomials. The results in [6] and [9] are based on hypothesis formulated using initial numerical experiments. Some rather surprising results have also been suggested to us by numerical experiments along the way. In particular, the relation between Faber polynomials and Chebyshev polynomials does not seem to have been given any attention in the literature, although it is known that they coincide for certain sets.
In short, we believe that the use of Tang’s algorithm in the study of Chebyshev polynomials may prove useful in the future when formulating well-posed hypothesis on their asymptotic behavior.
@@ Appendix A Tang’s algorithm
(33) $L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(g)=\sum_{j=1}^{n+1}r_{j}\mathrm{Re}(e^{-i\alpha_{j}}g(z_{j}))$
@@
(34) $A(\boldsymbol{z},\boldsymbol{\alpha})\coloneqq\begin{pmatrix}1&1&\cdots&1\\ \operatorname{Re}(e^{-i\alpha_{1}}\varphi_{1}(z_{1}))&\operatorname{Re}(e^{-i\alpha_{2}}\varphi_{1}(z_{2}))&\cdots&\operatorname{Re}(e^{-i\alpha_{n+1}}\varphi_{1}(z_{n+1}))\\ \vdots&\vdots&\ddots&\vdots\\ \operatorname{Re}(e^{-i\alpha_{1}}\varphi_{n}(z_{1}))&\operatorname{Re}(e^{-i\alpha_{2}}\varphi_{n}(z_{2}))&\cdots&\operatorname{Re}(e^{-i\alpha_{n+1}}\varphi_{n}(z_{n+1}))\end{pmatrix}$
@@
(35) $c_{f}(\boldsymbol{z},\boldsymbol{\alpha})\coloneqq\begin{pmatrix}\operatorname{Re}(e^{-i\alpha_{1}}f(z_{1}))\\ \vdots\\ \operatorname{Re}(e^{-i\alpha_{n+1}}f(z_{n+1}))\end{pmatrix}.$
@@
(36) $L_{\boldsymbol{r},\boldsymbol{\alpha},\boldsymbol{z}}(f)=c_{f}(\boldsymbol{z},\boldsymbol{\alpha})^{T}\boldsymbol{r}$
@@
(37) $A(\boldsymbol{z},\boldsymbol{\alpha})\boldsymbol{r}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}.$
Parameters $\boldsymbol{r},\,\boldsymbol{\alpha},\,\boldsymbol{z}$ satisfying (37) are called admisible if additionally $A(\boldsymbol{z},\boldsymbol{\alpha})$ is invertible. If $\varphi^{\ast}=\sum_{k=1}^{n}\lambda_{k}^{\ast}\varphi_{k}$, $\lambda_{k}^{\ast}\in\mathbb{R}$, is a best approximation and $\boldsymbol{r}^{\ast},$ $\boldsymbol{\alpha}^{\ast}$, $\boldsymbol{z}^{\ast}$ are corresponding admissible parameters such that $L_{\boldsymbol{r}^{\ast},\boldsymbol{\alpha}^{\ast},\boldsymbol{z}^{\ast}}(f)=\|f-\varphi^{\ast}\|_{\mathsf{E}}$ then
(38) $A(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})^{T}\begin{pmatrix}\|f-\varphi^{\ast}\|_{\mathsf{E}}\\ \lambda_{1}^{\ast}\\ \vdots\\ \lambda_{n}^{\ast}\end{pmatrix}=c_{f}(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$
and therefore if $A(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$ is invertible we can recover the extremal coefficients $\lambda_{1}^{\ast},\dotsc,\lambda_{n}^{\ast}$ from $A(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$ and $c_{f}(\boldsymbol{z}^{\ast},\boldsymbol{\alpha}^{\ast})$. We assume, as in [50, 51, 17] that $\mathsf{E}=I\eqqcolon[0,1]$ (which is no restriction) since we can always parametrize $\mathsf{E}$ using $[0,1]$. To emphasize that we are working on $[0,1]$ we let $\boldsymbol{z}=\boldsymbol{t}=\{t_{j}\}$.
An implementation of the algorithm in Python can be found at https://github.com/olguvirubin/complex_chebyshev
Input: Basis functions $\{\varphi_{k}\}_{k=1}^{n}$, $f\in\mathcal{C}_{\mathbb{R}}(I)$, $\{t_{j}^{(1)}\}_{j=1}^{n+1}$, $\{\alpha_{j}^{(1)}\}_{j=1}^{n+1},\mathrm{threshold}$
find: $\boldsymbol{r}$ satisfying $A(\boldsymbol{t}^{(1)},\boldsymbol{\alpha}^{(1)})\boldsymbol{r}=\begin{bmatrix}1\\ \boldsymbol{0}\end{bmatrix}$.
for $j\in\{1,\dotsc,n+1\}$ do
       if $r_{j}<0$ then
             let: $\alpha_{j}^{(1)}\coloneqq\alpha_{j}^{(1)}\pm\pi\in[0,2\pi)$
let: $\boldsymbol{r}^{(1)}\coloneqq A(\boldsymbol{t}^{(1)},\boldsymbol{\alpha}^{(1)})^{-1}\begin{bmatrix}1\\ \boldsymbol{0}\end{bmatrix}$.
  // This will produce admissible parameters
for $\nu=1,2,\cdots$ do
       compute: $\begin{bmatrix}h&\boldsymbol{\lambda}\end{bmatrix}^{T}=[A(\boldsymbol{t},\boldsymbol{\alpha})^{T}]^{-1}c_{f}(\boldsymbol{t},\boldsymbol{\alpha})$
@@
       let: $\varphi\coloneqq\sum_{k=1}^{n}\lambda_{k}\varphi_{k}$
      if $\|f-\varphi\|_{I}-h<\mathrm{threshold}\cdot h$ then
@@
             let: $\boldsymbol{d}$ be the solution to $\boldsymbol{d}\cdot A(\boldsymbol{t},\boldsymbol{\alpha})^{T}=\begin{bmatrix}1&\operatorname{Re}(e^{-i\vartheta}\varphi_{1}(x))&\cdots&\operatorname{Re}(e^{-i\vartheta}\varphi_{n}(x))\end{bmatrix}$
@@
             let: $\delta\coloneqq r_{\rho}/d_{\rho}$
             let: $t_{\rho}\coloneqq x$
             let: $\alpha_{\rho}\coloneqq\vartheta$
             for $j\in\{1\cdots n+1\}\setminus\{\rho\}$ do
                   $r_{j}\coloneqq r_{j}-\delta\cdot d_{j}$
            $r_{\rho}\coloneqq\delta$
@@
Algorithm 1 enables the computation of best approximations to complex valued functions.